OpenStudy (anonymous):

Emily uses her college photocopier machine to photocopy at the rate of $0.08 per page. She decides to rent a photocopier machine for $90 a year. The cost of photocopying using the rented machine is $0.03 per page. Part A: Write an inequality that can be used to calculate the number of pages that Emily should photocopy in a year so that the amount she pays for the rented machine is less than the college machine. Define the variable used. (5 points) Part B: How many pages should Emily print in a year to justify renting the photocopier? Show your work. (5 points)

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

using y = mx + b then cost of photocopying = 0.03 * number of pages + $90

OpenStudy (triciaal):

for rental

OpenStudy (anonymous):

May you show me which one is for part a and which one is for part b srry I get confused

OpenStudy (triciaal):

Emily uses her college photocopier machine to photocopy at the rate of $0.08 per page. cost of photocopying = 0.08 * number of pages

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

Part A how many pages to copy so that amount she pays for the rented machine is less than the college machine define the variables let the number of pages = p cost at school = C cost when rented = R find x when R = C the maximum number of pages

OpenStudy (anonymous):

thank you

OpenStudy (triciaal):

find when R < C

OpenStudy (anonymous):

?

OpenStudy (triciaal):

oops I also cover Part B in my answer above

OpenStudy (triciaal):

Part A She decides to rent a photocopier machine for $90 a year. 0.08 *x < 90 x =maximum number of pages she can print that would be the same cost of renting the machine

OpenStudy (triciaal):

do you get the general idea?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so that was part a right now I need part b please

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

How many pages should Emily print in a year to justify renting the photocopier? find when the cost are equal that's the number of pages see formula above

OpenStudy (anonymous):

may you do it please

OpenStudy (triciaal):

try using what I have above i'll work with you

OpenStudy (triciaal):

R = 0.03*x + 90 C = 0.08 * x when R = C what do you have?

OpenStudy (anonymous):

I have no clue

OpenStudy (triciaal):